OpenStudy (anonymous):

The function f(x) = (1.010979)^12x models the monthly interest that a bank offers to Dan after x years. Dan converts the function to have x isolated in the exponent. What is the approximate rate of growth? ANSWER CHOICES 1% 8% 12% 14%

OpenStudy (anonymous):

@inkyvoyd @nelsongreg6

OpenStudy (anonymous):

@BangkokGarrett

OpenStudy (anonymous):

1.010979 ^ 12 = 1.14 So you get 14 % f(x) = 1.14^x....again answer is 14%

OpenStudy (anonymous):

@BangkokGarrett you're awesome, can you help me with another one ??

OpenStudy (anonymous):

probably

OpenStudy (anonymous):

@BangkokGarrett heres a screenshot, i think its easier this way

OpenStudy (anonymous):

and just one more, i have the answer but i wanna confirm it for this... The function below shows the number of students of a school who enrolled for ballet classes. Let f(x) represent the total number of students who enrolled for the classes after x years. f(x) = 12(1.06)x The average rate of change in the number of students who enrolled for ballet classes from the third to the fifth year is ________students per year. Round your answer to the nearest whole number. the answer is 1 right

OpenStudy (anonymous):

2nd one....each value of f(n) is just the sum of the 2 prior values

OpenStudy (anonymous):

@BangkokGarrett ok awesome, can you confirm if my answer is right for that question i just posted really quick, its 1 right? after its rounded from 0.88 ...